# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers from Driving Video

### v2

@@ 2.3 Scenario-specific Video Sampling
The overall pipeline for scenario-specific video sampling is illustrated in Figure 4. Let $\mathcal{D}={D_{1},D_{2},\dots,D_{d}}$ be the set of subjects that were filtered out in section 2.2. Each subject $D_{i}$ performed a set of drives $dr_{1},dr_{2},\dots,dr_{n}$}} where $n$ is the maximum number of drives per subject. Each drive was further segmented into short video clips $V_{1},V_{2},\dots,V_{m}$}, where $m$ is the maximum number of video clips per drive. These video clips were recorded under different driving scenarios, $\mathcal{S}={Sc_{1},Sc_{2},\dots,Sc_{s}}$. From these video clips, a subset of scenario-specific video samples $V_{k}^{(D_{i},Sc_{j})}$ was filtered for each subject-scenario pair $(D_{i},Sc_{j})$ where $k$ is the index of the sampled video clip, $i$ is the index of the subject and $j$ is the index of the scenario. This resulted in a quality dataset of labeled video samples that were used as inputs to the large vision model in the next stage. This sampling strategy ensures sufficient data representation across both aging groups (Normal vs. MD) and allows the proposed method to be validated on balanced and identical driving scenarios. Additionally, the hierarchical video sampling steps across subjects, drives, and scenarios ensure a structured and scalable representation across varying cognitive profiles and driving contexts.
@@ 2.4 Large Vision Model-based Framework
The video samples derived in section 2.3 were fed into a pre-trained large Vision Transformer model [37] to generate vectorized representations of the video data. The objective of this framework is to encode driving videos into discriminative vector representations for downstream classification of cognitive aging groups. These representations capture prominent visual features that contrast the driving behavior of the two cognitive groups and distinguish them based on the vector’s distance in feature space (illustrated in Figure 5). The entire framework includes four steps as follows:
@@
$Z_{k}=F_{\theta}(V_{k}^{(D_{i},Sc_{j})}),\quad Z_{k}\in\mathbb{R}^{m}$
@@
$\tilde{Z}_{k}=g(Z_{k}),\quad\tilde{Z}_{i}\in\mathbb{R}^{n},\text{ where }n\in\{50,100,200\},n\ll m$
@@
To evaluate the relevance of a scenario $Sc_{j}\in\mathcal{S}$ for distinguishing the two aging groups, we computed a class-wise Euclidean distance in the reduced feature space. In Figure 5, an example of a dimension-reduced representation ($n=2$) with the group separability in feature space is illustrated. The Euclidean distance between the class centroids was obtained as follows:
$\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{normal}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{normal}}^{(Sc_{j})}}\mathbf{z},\quad\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{MD}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{MD}}^{(Sc_{j})}}\mathbf{z}$
@@
$\mathrm{Dist}(Sc_{j})=\left\|\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}-\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\right\|_{2}$
@@
$Sc_{\max}=\arg\max_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$
@@
$Sc_{\min}=\arg\min_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$
Let $\tilde{\mathcal{Z}}^{(Sc)}=\{\tilde{Z}_{k}^{(Sc)}\}$ denotes the reduced dimentional vectors (from Step 2) and $\mathcal{Y}^{(Sc)}=\{Y_{k}^{(Sc)}\}$ their corresponding labels for scenario $Sc\in\{Sc_{\max},Sc_{\min}\}$. We trained two RF classifiers with identical hyperparameters on the reduced feature space for both scenarios $\{Sc_{\max},Sc_{\min}\}$:
$\hat{Y}_{k}^{(Sc)}\;=\;\tilde{H}^{(Sc)}\!\left(\tilde{Z}_{k}^{(Sc)}\right),\qquad\tilde{H}^{(Sc)}:\mathbb{R}^{n}\rightarrow\mathcal{Y},$
@@
$\bigl{\{}\mathbf{a},\,\mathbf{P},\,\mathbf{R},\,\mathbf{F1}\bigr{\}}_{Sc_{\max}}\quad\text{vs.}\quad\bigl{\{}\mathbf{a},\,\mathbf{P},\,\mathbf{R},\,\mathbf{F1}\bigr{\}}_{Sc_{\min}}$
and reported the performance difference as a validation criterion. A consistent performance gain of $\tilde{H}^{(Sc_{\max})}$ over $\tilde{H}^{(Sc_{\min})}$ validates the proposed scenario-based approach by demonstrating that scenarios with larger inter-centroid separation yield stronger discriminability between the two cognitive aging groups. The entire framework is shown in Algorithm 1. Moreover, we further evaluated the validity of the framework under different data sampling strategies. Multiple sampling strategies (random and driver-level separation) were employed to assess whether the observed RF classifier performance generalizes across subjects and settings. We reported the average performance $\mathbf{a,P,R,}$ and $\mathbf{F1}$ under two complementary sampling strategies:
@@
12:  Divide $\tilde{\mathcal{E}}_{Sc_{j}}$ into class-wise sets $\tilde{Z}_{\text{normal}}^{(Sc_{j})},\tilde{Z}_{\text{MD}}^{(Sc_{j})}$
13:  $\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}\!\leftarrow\!\frac{1}{|\tilde{Z}_{\text{normal}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{normal}}^{(Sc_{j})}}\mathbf{z}$, $\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\!\leftarrow\!\frac{1}{|\tilde{Z}_{\text{MD}}^{(Sc_{j})}|}\sum_{\mathbf{z}\in\tilde{Z}_{\text{MD}}^{(Sc_{j})}}\mathbf{z}$
@@
21:   Stratified train/val/test split on samples (balanced)
@@
23:   Leave-$k$-drivers-out: hold out $k$ drivers as test set; repeat $r$ runs with different held-out drivers
24:  end if
25:  Train RF $\tilde{H}^{(Sc)}:\mathbb{R}^{n}\to\mathcal{Y}$ on train split $\triangleright$ Step 4
26:  Evaluate on held-out data to obtain $\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc}$; average over runs if DLS
27:end for
28:return $\{\mathrm{Dist}(Sc_{j})\}_{Sc_{j}\in\mathcal{S}},\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc_{\max}},\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc_{\min}}$
@@
This sampling strategy evaluates both intra-subject consistency (random) and inter-subject generalization (DLS) in the proposed modeling approach.
@@ 3.1 Scenario-based Roadway Selection
To investigate the connection between driving behavior and cognitive status, we tested and selected two distinct scenarios from the Omaha most transited route where older drivers exhibited the most discriminative driving patterns. These patterns can serve as interpretable behavioral cues for identifying cognitive decline. The selected driving scenarios are:
@@
    Interstate: high-speed, lane-keeping scenarios on long uninterrupted roads. e.g., Omaha I-80 W, I-80 E.
(a) Interstate segments
(b) Freeway-interchange segments
Figure 6: GIS map - (a) Interstate and (b) Freeway-interchange scenarios. The relevant route segments were filtered out from the Omaha most transited route database shown in Figure 3.
From different federal functional categories (e.g., Local, Major collector, Minor collector, etc.), we selected the most contrastive scenario combination and performed scenario-specific video sampling as section 2.3. Figure 6 shows the GIS map highlighting the two driving scenarios for our analysis. We considered a single drive having $\leq 0.31$ miles or $\leq 500$ meters in length as a route segment. Omaha most transited route database consists of 24,106 route segments. In total, we filtered out $964$ unique route segments, consisting of $632$ freeway-interchange segments (highlighted in green on Figure 6(b)) and $332$ interstate segments (highlighted in blue on Figure 6(a)). Table 2 summarizes the key statistics – a balance of the selected data and driving exposure of the cognitive groups. Additionally, we selected the route segments that cover at least a minimum number of drives for both cognitive groups. To interpret the driving exposure data distribution in Table 2, we defined several key variables below:
@@
Table 2: Driving exposure in the route segments for two scenarios
Scenario Normal MCI MD Unique Recurring Recurring trip
Fwy-int. 8136 10134 919 19189 68048 3.55
Interstate 7123 7894 780 15797 69620 4.41
We leveraged “unique drives” vs. “repeated drives” per subject across all the route segments to indicate statistical reliability and diversity in driving exposure by the two cognitive groups. The recurring trips per subject suggest that each participant had substantial driving trips with the selected roadway scenario, offering a significant amount of video data for our analysis.
Figure 7 shows a boxplot distribution of speed limits and the trip distance for both scenarios. As we observed, the freeway-interchange segments tend to have lower speed limits and shorter distances, which work as a transitional driving scenario, involving challenging driving maneuvers and fast decision-making. In contrast, the interstate segments exhibit higher speed limits and longer travel distances, which indicate an uninterrupted highway driving scenario with relaxed driving behavior.
@@
These characteristics offer a strong contrast between the two scenarios, reflecting distinct driving demands and cognitive load associated with each scenario.
3.2 Executing Video-based Model
We conducted the section 2.4 experiments across the two representative scenarios with both transitional and straight driving maneuvers. The aim was to investigate whether the large vision model could capture the subtle visual features to effectively distinguish driving behavior relevant for cognitive assessment. The experiments were conducted under the following sampling protocols (step 4 in section 2.4):
@@ 3.2 Executing Video-based Model
    Random Sampling: Video segments are randomly split into train and test sets without enforcing driver-level separation.
@@
In our experiments, the parameters were set as: (1) Input image size (960x752x3), (2) Vision model backbone: ResNet-v2 [41] (3) Video-to-vector dimension (6144x1), (4) Frame rate $\{1,10\}$ Hz, (5) Low-dimension size, $n\in\{50,100,200\}$, and (6) Distance metric: Euclidean distance. This setting provided a comprehensive understanding of how video-derived feature spaces behave under varying sampling conditions and data splits and laid the foundation for downstream classification tasks.
@@ 4.1 Model Performance on Different Scenarios
Table 3: Model performance on two roadway types
@@
As discussed in section 2.4, we trained two RF models to identify the aging-groups (Normal-aging vs. MD-aging) across the Freeway interchange and Interstate scenarios. Table 3 reveals the performance of the trained RF models. The $\Delta$ row (without driver-level separation) indicates significant separation on all metrics $\mathbf{(a,P,R,F1)}$ and validates the selection of these contrastive scenarios to classify aging groups. The performance difference (grey rows) between the two scenarios indicates that there are prominent vision features in these specific scenarios that can successfully distinguish the aging groups solely by looking at the driving behavior features.
Furthermore, the driver-level separation yields slightly widened separation, which can be observed on all metrics $\mathbf{(a,P,R,F1)}$ in the $\Delta_{\text{ds}}$ row. It surpassed the performance difference achieved without driver-level separation; for example, the accuracy difference went up from $\Delta$ = 15.91% to $\Delta_{\text{ds}}$ = 17.64%, indicating a 1.73% increment in performance difference. The $\Delta_{\text{ds}}$ aligns with achieving a significant performance difference while there is no “subject-level overlapping”. It also validates that the Fwy-int scenario includes more prominent driving movement features compared to its counterpart. On the other hand, the Interstate scenario performs significantly worse, yielding lower accuracy and F1-score. Therefore, the Freeway interchange is a more behaviorally informative and cognitively sensitive driving scenario, where complex maneuvers, short distances and lower speed variability have more nuanced driver responses.
We compared our results with a recent study [42] that used a multi-modal (demographics and sleep log data) framework to detect cognitive decline in older drivers. Our method outperformed both without Driver-level Separation ([42] accuracy 68.64% vs. our Fwy-int. accuracy 71.03%) and with Driver-level Separation ([42] accuracy 70.48% vs. our Fwy-int. max accuracy 72.32%) cases on the RWRAD dataset. This highlights the effectiveness of scenario-specific vision features as non-invasive digital biomarkers for cognitive aging classification. Moreover, another recent study by Danda et al. [43] found that amyloid-positive older adults exhibited higher average speeds, intermittent speed variability and increased acceleration on freeway on-ramp driving (similar to Fwy-int scenario), suggesting less adaptive driving. In clinical research, Amyloid is a protein that accumulates abnormally in the brain and forms plaques linked to the development of MD. Patients with cognitive impairment who test amyloid-positive are typically diagnosed as having MD, whereas amyloid-negative patients are generally excluded from an MD diagnosis [44, 45]. These findings confirm that Freeway ramp driving requires increased cognitive load involving behavioral reactions to rapidly changing traffic and abrupt speed shifts, combining perception with motor control [43]. Additionally, they underscore our results of higher classification accuracy ($\mathbf{a}$ = 71.03%) and F1-score ($\mathbf{F1}$ = 0.6712) in Freeway interchange scenarios compared to the Interstate settings ($\mathbf{a}$ = 55.12%, $\mathbf{F1}$ = 0.5368). Although their approach combined driving patterns with physiological signals, it validates the relevance of using contrastive and challenging roadway scenarios in cognitive assessment tasks. Therefore, this study supports our findings that the Freeway interchange scenario provides more informative and strong behavioral cues that the large vision model can leverage to better separate cognitive aging groups.

### v3

@@ 2.3 Scenario-specific Video Sampling
The overall pipeline for scenario-specific video sampling is illustrated in Fig. 5. Let $\mathcal{D}={D_{1},D_{2},\dots,D_{d}}$ be the set of subjects that were filtered out in section 2.2. Each subject $D_{i}$ performed a set of drives $dr_{1},dr_{2},\dots,dr_{n}$}} where $n$ is the maximum number of drives per subject. Each drive was further segmented into short video clips $V_{1},V_{2},\dots,V_{m}$}, where $m$ is the maximum number of video clips per drive. These video clips were recorded under various driving scenarios, $\mathcal{S}={Sc_{1},Sc_{2},\dots,Sc_{s}}$. From these video clips, a subset of scenario-specific video samples $V_{k}^{(D_{i},Sc_{j})}$ was filtered for each subject-scenario pair $(D_{i},Sc_{j})$ where $k$ is the index of the sampled video clip, $i$ is the index of the subject and $j$ is the index of the scenario. This resulted in a quality dataset of labeled video samples that were used as inputs to the large vision model in the next stage. This sampling strategy ensures sufficient data representation across both aging groups (Normal vs. MD) and allows the proposed method to be validated on balanced and identical driving scenarios. Additionally, the hierarchical video sampling steps across subjects, drives, and scenarios ensure a structured and scalable representation across varying cognitive profiles and driving contexts.
@@ 2.4 Large Vision Model-based Framework
The video samples derived in section 2.3 were fed into a pre-trained large Vision Transformer model [37] to generate vectorized representations of the video data. The objective of this framework is to encode driving videos into discriminative vector representations for downstream classification of cognitive aging groups. These representations capture prominent visual features that contrast the driving behavior of the two cognitive groups and distinguish them based on the vector’s distance in feature space (illustrated in Fig. 6). The entire scenario-based video analysis framework is shown in Algorithm 1 which includes the following four steps:
@@
$Z_{k}=F_{\theta}(V_{k}^{(D_{i},Sc_{j})}),\quad Z_{k}\in\mathbb{R}^{m}$ (1)
@@
$\tilde{Z}_{k}=g(Z_{k}),\quad\tilde{Z}_{i}\in\mathbb{R}^{n},\quad n\in\{50,100,200\},\;n\ll m$ (2)
@@
To evaluate the relevance of a scenario $Sc_{j}\in\mathcal{S}$ for distinguishing the two aging groups, we computed a class-wise Euclidean distance in the reduced feature space. In Fig. 6, an example of a dimension-reduced representation ($n=2$) with the group separability in feature space is illustrated. The Euclidean distance between the class centroids was obtained as follows:
$\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{normal}}^{(Sc_{j})}|}\sum\limits_{\mathbf{z}\in\tilde{Z}_{\text{normal}}^{(Sc_{j})}}\mathbf{z}$ (3)
$\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}=\frac{1}{|\tilde{Z}_{\text{MD}}^{(Sc_{j})}|}\sum\limits_{\mathbf{z}\in\tilde{Z}_{\text{MD}}^{(Sc_{j})}}\mathbf{z}$ (4)
@@
$\mathrm{Dist}(Sc_{j})=\left\|\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}-\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\right\|_{2}$ (5)
@@
$Sc_{\max}=\arg\max_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$ (6)
@@
$Sc_{\min}=\arg\min_{Sc_{j}\in\mathcal{S}}\;\mathrm{Dist}(Sc_{j})$ (7)
@@
12:  Divide $\tilde{\mathcal{E}}_{Sc_{j}}$ into two class-wise sets: $\mathbf{z_{n}}=\tilde{Z}_{\text{normal}}^{(Sc_{j})}$,   $\mathbf{z_{m}}=\tilde{Z}_{\text{MD}}^{(Sc_{j})}$
13:  $\boldsymbol{\mu}_{\text{normal}}^{(Sc_{j})}\leftarrow\frac{1}{|\mathbf{z_{n}}|}\sum\limits_{\mathbf{z}\in\mathbf{z_{n}}}\mathbf{z}$,   $\boldsymbol{\mu}_{\text{MD}}^{(Sc_{j})}\leftarrow\frac{1}{|\mathbf{z_{m}}|}\sum\limits_{\mathbf{z}\in\mathbf{z_{m}}}\mathbf{z}$
@@
21:    train/val/test split on samples (balanced)
@@
23:    Leave-$k$-drivers-out: hold out $k$ drivers as test set;
24:    Repeat $r$ runs with different held-out drivers
25:  end if
26:  Train RF $\tilde{H}^{(Sc)}:\mathbb{R}^{n}\to\mathcal{Y}$ on train split $\triangleright$ Step 4
27:  Evaluate on held-out data to obtain $\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc}$;
28:  Average over runs if DLS
29:end for
30:return $\{\mathrm{Dist}(Sc_{j})\}_{Sc_{j}\in\mathcal{S}},\{\mathbf{a},\mathbf{P},\mathbf{R},\mathbf{F1}\}_{Sc_{j}\in\{Sc_{\max},Sc_{\min}\}}$
A consistent performance gain of $\tilde{H}^{(Sc_{\max})}$ over $\tilde{H}^{(Sc_{\min})}$ validates the proposed scenario-based approach by demonstrating that scenarios with larger inter-centroid separation yield stronger discriminability or contrast between the two cognitive aging groups. Moreover, we further evaluated the validity of the framework under different data sampling strategies. Multiple sampling strategies (random and driver-level separation) were employed to assess whether the observed RF classifier performance generalizes across subjects and settings. We reported the average performance $\mathbf{a,P,R,}$ and $\mathbf{F1}$ under two complementary sampling strategies:
@@
This sampling strategy evaluates both intra-subject consistency (random) and inter-subject generalization (DLS for unseen subjects) in the proposed modeling approach.
@@ 3.1 Scenario-based Roadway Selection
To investigate the connection between driving behavior and cognitive load under different driving scenarios, we tested and selected two scenarios from the OMR data where older drivers exhibited the most discriminative driving patterns. These patterns can serve as interpretable behavioral cues for identifying cognitive decline. The selected driving scenarios are:
(a) Interstate segments
(b) Freeway-interchange segments
Figure 7: GIS map for (a) Interstate and (b) Freeway-interchange scenarios. The relevant route segments were filtered out from the OMR database shown in Fig. 4.
@@
    Interstate: high-speed, lane-keeping scenarios on long and uninterrupted roads. e.g., Omaha I-80 W, I-80 E.
From different federal functional categories (e.g., Local, Major collector, Minor collector, etc.), we selected the most contrastive scenario combination and performed scenario-specific video sampling as section 2.3. Fig. 7 shows the GIS map highlighting the two driving scenarios for our analysis. We considered a single drive with a length of $\leq 0.31$ miles or $\leq 500$ meters as a route segment. OMR database consists of 24,106 route segments. In total, we filtered out $964$ unique route segments, consisting of $632$ freeway-interchange segments (highlighted in green on Fig. 7(b)) and $332$ interstate segments (highlighted in blue on Fig. 7(a)). Table 2 summarizes the key statistics – a balance of the selected data and driving exposure of the cognitive groups. Additionally, we selected the route segments that cover at least a minimum number of drives for both cognitive groups. To interpret the driving exposure data distribution in Table 2, we defined several key variables below:
Table 2: Driving exposure in the OMR route segments
Scenario Normal MCI MD Unique Recurring Recurring trip
Fwy-int. 8136 10134 919 19189 68048 3.55
Interstate 7123 7894 780 15797 69620 4.41
@@
We leveraged “unique drives” vs. “repeated drives” per subject across all route segments to indicate statistical reliability and diversity in driving exposure by the two cognitive groups. The recurring trips per subject suggest that each participant had substantial driving trips with the selected roadway scenario which offers a significant amount of representative video data for our analysis.
Fig. 8 shows a boxplot distribution of speed limits and the trip distance for both scenarios. As we observed, the freeway-interchange segments tend to have lower speed limits and shorter distances, which work as a transitional driving scenario, involving challenging driving maneuvers and fast decision-making. In contrast, the interstate segments have higher speed limits and longer travel distances, which indicate an uninterrupted highway driving scenario with relaxed driving behavior.
@@
These characteristics offer a strong contrast between the two scenarios. In addition, they indicate distinct driving behaviors and cognitive load associated with each scenario. As suggested in [16], freeway merging requires substantially higher cognitive demand than straight driving due to rapid integration of sensory processing, decision-making and cognitive responses to adapt to fast-moving traffic and sudden speed changes.
3.2 Implementation of Video-based Model
We conducted the experiments across two representative scenarios with both transitional and straight driving maneuvers. The aim was to investigate whether the large vision model could capture the subtle visual features to effectively distinguish driving behavior relevant for cognitive assessment. The experiments were conducted under the following sampling protocols (step 4 in section 2.4):
@@ 3.2 Implementation of Video-based Model
    Random Sampling: Video segments were randomly split into train and test sets without enforcing any driver-level separation.
@@
[…]
@@ 4.1 Model Performance on Different Scenarios
Table 3: Model performance on the OMR subset for two contrastive scenarios
@@
As discussed in section 2.4, we trained two RF classifier models to classify the aging-groups (Normal-aging vs. MD-aging) across the Freeway interchange and Interstate scenarios of the OMR data. Table 3 shows the performance of the trained RF models. The results are reported under both random sampling and driver-level separation (DLS) protocols, with the grey rows showing the performance differences ($\Delta$ and $\Delta_{\text{ds}}$).
As observed, the Freeway interchange scenario consistently outperforms the Interstate scenario across all metrics $\mathbf{(a,P,R,F1)}$, achieving 71.03% accuracy and an $\mathbf{F1}$ score of 0.7662 under random sampling. The performance gaps, highlighted in the grey rows ($\Delta$ and $\Delta_{\text{ds}}$), confirm that certain roadway scenarios contain more discriminative vision features for distinguishing Normal-aging and MD-aging drivers. Furthermore, including the driver-level separation (DLS) yields an increase in performance difference, which can be observed on all metrics in the $\Delta_{\text{ds}}$ row. It surpassed the performance difference achieved without the DLS case; for example, the accuracy difference went up from $\Delta$ = 15.91% to $\Delta_{\text{ds}}$ = 17.64%, indicating a 1.73% increment in performance difference. Therefore, the Freeway interchange is a more behaviorally informative and cognitively sensitive driving scenario, where complex maneuvers, short distances and lower speed variability have more nuanced driver responses.